OpenStudy (anonymous):

Yesterday, Jed's bank account balance was –$26.34. He deposited some money into his account so that the balance is now $42.18. How much did he deposit into his account? A. $15.84 B. $16.84 C. $68.42 D. $68.52

OpenStudy (anonymous):

42.18 - (-26.34) =15.84

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

what do i do with the middle

OpenStudy (anonymous):

of 42.18 - (-26.34) = 42.18 + 26.34 = ?

OpenStudy (anonymous):

the =

OpenStudy (bishoppatton):

42.18 + 26.34 = 68.52

OpenStudy (anonymous):

so D?

OpenStudy (bishoppatton):

Yes.

OpenStudy (anonymous):

thanks :)

OpenStudy (bishoppatton):

Np

OpenStudy (bishoppatton):

what did you get?